Two roda A and B of same length and radius are joined together. The thermal conductivity of A and B are 2K and K. Under steady state conditions ,if the temperature difference between the open ends of A and B is 36°C , the temperature difference across `A` is

Grade:12

1 Answers

Shailendra Kumar Sharma
6 years ago
Q=\frac{T2-T1}{L/KA+L/2KA} 
Where T2-T1 =36 
Now accross A
Q=\frac{T2-T1}{L/2KA}
Where T2-T1 = TA (say)Q=\frac{TA}{L/2KA}=\frac{36}{L/2KA+L/KA}
Q=\frac{T2-T1}{L/2KA}=\frac{T2-T1}{L/2KA+L/KA}
Q=\frac{TA}{L/2KA}=\frac{36}{3L/2KA}
TA =12 so the Temperature difference will be 12 degree